# Reducing Hallucinations in Language Model-based SPARQL Query Generation Using Post-Generation Memory Retrieval

### v2

@@ 4.1.3 Setup
finetuned 74.23 % 92.76 80.15 % 0.38 %
Llama 3.1 8B finetuned 0.72 % 64.83 0.79 % 89.84 %
@@
few-shot 43.53 % 73.53 69.60 % 0.0 %
Table 1: Results on LCQUAD 2.0 (original split): Across different LLMs, PGMR demonstrates a substantial reduction in hallucinations while achieving significantly higher Query EM scores than direct SPARQL generation, as discussed in Section 6.
@@ 5.1 Datasets
We evaluate our work on two widely used datasets for SPARQL query generation: LCQUAD 2.0 (Dubey et al., 2019) and QALD-10 (Usbeck et al., 2023). Even though tagged versions of these datasets exist, where the entities and relations in the question are already linked and tagged with associated URIs from the KG, we use the untagged versions. Tagging is expensive, time-consuming, and impractical in the real world. SPARQL query generation becomes a much harder problem when using untagged data as it forces the language models to generate KG elements like URIs based on their internal parametric knowledge, which leads to hallucinations and out-of-distribution errors, as discussed in Section 6.
LCQUAD 2.0 comprises a mix of simple and complex questions with associated SPARQL queries over Wikidata, crafted by human annotators from Amazon Mechanical Turk. The original split of LCQUAD 2.0 includes a broad dataset with 21k questions for training, 3k questions for validation, and 6k questions for testing. To specifically test how our models perform in the face of out-of-distribution URIs not seen before in training, we also construct an "unknown URI" split of LCQUAD 2.0 in line with Reyd and Zouaq (2023). Every query in the test set of this version of the dataset includes at least one URI not seen during training, thus making accurate query generation a much harder problem. The "unknown URI" split contains 24k train, 3k validation, and 3k test questions and associated queries.
QALD-10 is a benchmarking dataset that is a part of the Question Answering over Linked Data (QALD) challenge series for KGQA using SPARQL query generation. Owing to the complexity of SPARQL queries that involve various clauses and literals, QALD-10 is recognized as one of the most challenging and practically applicable datasets in the QALD challenge series. Even though each question in QALD-10 is translated into 8 different languages, we only consider the English versions of the questions. The dataset consists of 412 training questions and 394 test questions, which is much smaller than LCQUAD 2.0 but contains more complex queries. Given the limited availability of training data for QALD-10, we employ a two-step approach in our finetuning experiments: pre-training the LLM on LCQUAD 2.0 before finetuning them on QALD-10.
@@ 5.2 Metrics
Query EM: We evaluate the query exact match score (Query EM), which directly compares if the predicted query and the reference query are an exact match.
URI EM: We assess the URI exact match (URI EM) score, a metric that directly evaluates whether the set of all URIs in the predicted SPARQL query matches exactly with those in the reference query.
@@
Model Type Query EM $\uparrow$ BLEU $\uparrow$ URI EM $\uparrow$ URI Hallucination $\downarrow$
T5-Small finetuned 0.16 % 62.91 0.36 % 61.66 %
T5-Small PGMR
finetuned 59.97 % 86.28 78.06 % 4.63 %
Llama 3.1 8B finetuned 0.03 % 65.30 0.06 % 91.01 %
Llama 3.1 8B PGMR
finetuned 71.02% 89.62 79.49 % 0.0 %
GPT 4o few-shot 0.89 % 52.02 2.18 % 86.17 %
GPT 4o PGMR
few-shot 55.61 % 83.41 68.08 % 0.0 %
Table 2: Results on LCQUAD 2.0 (unknown URI split): Even in an out-of-distribution setting, PGMR significantly mitigates hallucinations and enhances Query EM performance compared to direct SPARQL generation across various LLMs, as discussed in Section 6.
@@ 5.3 Baselines
PGMR: As described in Section 4.1, we implement the proposed PGMR, where SPARQL queries are first converted into an intermediate query, followed by finetuning or few-shot prompting an LLM on this intermediate representation. Our retriever then utilizes a non-parametric memory module to translate the intermediate queries generated by the LLM back into SPARQL queries.
@@ 5.3.1 Language Models
T5: Despite the development of various models for generating SPARQL queries with tagged questions (Banerjee et al., 2022; Diallo et al., 2024; Qi et al., 2024), finetuned T5 (Raffel et al., 2020) is currently the state-of-the-art model for generating SPARQL queries from untagged data on LC-QUAD 2 (Diallo et al., 2024). We use the "t5-small" model in our study, as Banerjee et al. (2022) show a limited benefit of the added parameters when using a larger T5 model for SPARQL query generation.
Llama 3.1 8B: We finetune Llama 3.1 8B Instruct (Dubey et al., 2024) to evaluate and compare direct generation to PGMR in our experiments.
GPT 4o: Additionally, we employ 25-shot prompting with GPT 4o (Achiam et al., 2023) to assess and compare the few-shot performance of direct generation with PGMR.
Unlike the evaluation approach utilized by Banerjee et al. (2022) and Qi et al. (2024), which involves generating multiple queries for each question and selecting the first query that successfully executes and returns an answer as the predicted query, our evaluation generates only a single query from our models. Additionally, since LLMs can often generate identical queries with variable names (e.g., ?$value$ and ?$s$) different from those found in the gold query (e.g., ?$uri$ and ?$sbj$), we normalize the variable names (e.g., ?$var0$ and ?$var1$) in our generated SPARQL queries in line with the evaluation approach followed by Qi et al. (2024).
@@ 6 Results
PGMR drastically reduces hallucinations. One of the key limitations of LLMs is their tendency to “hallucinate,” generating content that appears valid but lacks factual accuracy. This problem is especially pronounced in LLM-based direct SPARQL query generation, where LLMs often fabricate URIs that are not present in the knowledge graph. As seen in Table 1, PGMR reduces URI hallucinations compared to direct SPARQL generation by 27.37% for T5-small, 89.83% for Llama 3.1 8B, and 81.81 % for GPT 4o. This improvement is even more substantial on the unknown URI split of LCQUAD 2.0, where PGMR reduces URI hallucinations by 57.03% for T5-small, 91.01% for Llama 3.1 8B, and 86.17% for GPT 4o, as demonstrated in Table 2. On the harder QALD-10 dataset, direct SPARQL query generation increases URI hallucinations to 66.50% for T5-small, 96.14% for Llama 3.1 8B, and 83.76% for GPT 4o, while PGMR completely eliminates hallucinations for all LLMs tested, as seen in Table 3. These results highlight the effectiveness of PGMR in mitigating URI hallucinations regardless of the LLM, dataset, and data distribution used, with most cases demonstrating an almost complete suppression of hallucinations. This improvement is largely driven by PGMR’s approach, which utilizes a non-parametric memory module for URI retrieval following the LLM-based generation of an intermediate SPARQL-formatted query, bypassing the dependence on the LLM’s internal parametric knowledge.
Model Type Query EM $\uparrow$ BLEU $\uparrow$ URI EM $\uparrow$ URI Hallucination $\downarrow$
T5-Small finetuned 5.83 % 25.49 6.85 % 66.50 %
T5-Small PGMR
finetuned 18.27 % 30.26 20.30 % 0.0 %
Llama 3.1 8B finetuned 0.25 % 21.74 0.51 % 96.14 %
Llama 3.1 8B PGMR
finetuned 17.22 % 29.40 27.76 % 0.0 %
GPT 4o few-shot 3.30 % 38.63 7.10 % 83.76 %
GPT 4o PGMR
few-shot 25.52 % 51.66 44.33 % 0.0 %
Table 3: Results on QALD-10: PGMR consistently minimizes hallucinations across LLMs while attaining significantly better Query EM scores than direct SPARQL generation, even for the more complex QALD-10 dataset, as discussed in Section 6.
PGMR enhances the quality of SPARQL queries. Across all assessed LLMs, PGMR consistently enhances the accuracy and structure of SPARQL queries on the untagged versions of the LCQUAD 2.0 and QALD-10 datasets. As seen in Table 1, the application of our PGMR approach yields a remarkable enhancement in query EM, achieving a boost of nearly 37% for T5, 69% for Llama 3.1 8B, and 42% for GPT 4o over the traditional direct SPARQL query generation on the original split of LCQUAD 2.0. PGMR also improves the average BLEU score by almost 21 points across LLMs on the original split of LCQUAD 2.0. As seen in Table 2, this improvement is particularly pronounced on the out-of-distribution unknown URI split of LCQUAD 2.0, where PGMR achieves an average query EM score of more than 62% across LLMs tested, significantly outperforming direct generation, which attains only 0.36% on average. This further highlights the challenges faced by LLMs in directly generating SPARQL queries when the test data is out-of-distribution, and some of the URIs have not been seen during training. On the QALD-10 dataset, our PGMR approach enhances query EM performance over direct generation by approximately 13% for T5, 17% for Llama 3.1 8B, and 22% for GPT 4o, as seen in Table 3. Our results confirm that PGMR not only minimizes hallucinations to a great extent but also improves the accuracy and reliability of SPARQL query generation across various LLMs, datasets, and data distributions.
PGMR shows resilience to data distribution shifts. PGMR proves to be more resilient to distributional changes than direct SPARQL query generation, demonstrating superior generalization across LLMs as the dataset transitions from the original split of LCQUAD 2.0 to the unknown URI split (as discussed in Section 5.1). As shown in Tables 1 and 2, PGMR mitigates out-of-distribution performance degradation, with its average query EM score across LLMs remaining nearly unchanged, decreasing marginally from 62.4% to 62.2% when transitioning from the original split to the unknown URI split of LCQUAD 2.0. This is in sharp contrast to direct query generation, which experiences a substantial degradation in performance, as its query EM score drops from nearly 12% to a mere 0.36% on average across LLMs. This superior out-of-distribution performance stems from the fundamental difference between PGMR and traditional direct SPARQL query generation. While the latter depends on the LLM’s internal parametric knowledge for URI generation, PGMR instead leverages a non-parametric memory module to handle KG elements, allowing the LLM to focus on generating syntactically correct SPARQL queries.
PGMR increases URI EM. PGMR exhibits a substantial enhancement in URI exact match (EM) accuracy across a diverse range of LLMs, datasets, and data distributions, further establishing its advantages over conventional direct query generation methods, as evidenced by Tables 1, 2, and 3. On the original LCQUAD 2.0 split, PGMR substantially improves the accuracy of URIs in the generated SPARQL queries and achieves significant URI EM improvements, with score increases of 41.71% for T5, 77.97% for Llama 3.1 8B, and 65.24% for GPT-4o over conventional direct query generation techniques. The impact of PGMR is even more pronounced when the data distribution shifts to the unknown URI split of LCQUAD 2.0, with PGMR gaining 77.7% for T5, 79.43% for Llama 3.1 8B, and 65.9% for GPT 4o over the direct query generation approach, which experiences significant performance degradation across all evaluated LLMs. This trend continues on QALD-10, where PGMR improves the URI EM accuracy by 13.45% for T5, 27.25% for Llama 3.1 8B, and 37.23% for GPT 4o over direct SPARQL query generation.

### v3

@@ 4 Post-Generation Memory Retrieval (PGMR)
FT 86.75 86.37 % 94.63 93.17 % 97.05 % 0.0 %
Llama 3.1 8B PEFT 19.89 0.75 % 66.74 1.52 % 61.23 % 89.84 %
Llama 3.1 8B RAG PEFT 81.66 81.37 % 94.65 90.19 % 91.83 % 7.23 %
@@
25-shot 62.23 51.24 % 77.49 79.12 % 68.10 % 0.01 %
Table 1: Results on LCQUAD 2.0 (original split): Across different LLMs, PGMR demonstrates a substantial reduction in hallucinations while achieving significantly higher SQM and answer-level F1 scores than direct SPARQL generation and RAG, as discussed in Section 6.
4.1 Training Data Transformation
To train models capable of generating SPARQL-PGMR queries, we design a systematic transformation process that converts standard question-SPARQL query pairs into the PGMR representation. The transformation consists of three stages: entity replacement, relation replacement, and mapping construction. In the first stage, all entity URIs (e.g., Q937) appearing in the query are replaced with symbolic placeholders of the form entityX. A mapping entry is then created for each placeholder, expressed as entityX = [ENT] Label [/ENT] Description, where the label corresponds to the canonical surface form and the description provides a short disambiguating definition extracted from the KG (e.g., [ENT] Albert Einstein [/ENT] German-born theoretical physicist, developer of the theory of relativity, Nobel Prize laureate (1921)). The second stage applies the same procedure to relation and property URIs (e.g., P35), replacing them with placeholders such as relationY and constructing mappings of the form relationY = [REL] Label [/REL] Description. The description obtained from the KG provides the necessary semantic information for relation disambiguation (e.g., [REL] head of state [/REL] official with the highest formal authority in a country/state). Finally, a mapping block is appended to each transformed query, listing all entity and relation definitions. In this study, we obtain labels and descriptions from Wikidata using its official Python library [^4] https://pypi.org/project/Wikidata/.
4.2 Memory Retriever
Memories: We utilize two memory modules (one for entities and another for relations) designed as a dictionary that pairs LLM encoder-generated embeddings of URI labels and their descriptions as keys with the corresponding URIs as values. The entity memory module includes such embeddings and URIs for all entity URIs in the dataset. Similarly, the ontology memory module includes these for all the relation URIs in the dataset.
Retriever: Given a natural language input such as "[ENT] Albert Einstein [/ENT] German-born theoretical physicist, developer of the theory of relativity, Nobel Prize laureate (1921)," our memory retriever locates the URI whose label and description are most similar to this text within the relevant memory. The label is first parsed and matched with the URI labels in the memory. If there is no direct match or multiple matches for the label, then the input text with the label and description is encoded into an embedding by an LLM encoder. The retriever then uses FAISS (Douze et al., 2024; Johnson et al., 2019) to retrieve the URI from the memory whose embedding is nearest to the input text embedding. We use BGE (Chen et al., 2024) as the encoder for our retriever in this study.
4.3 Setup
Training Setup: Using the transformed training data from Section 4.1, we finetune the LLM to generate intermediate queries given the questions.
Few-shot Setup: Drawing on the transformed training data, we apply few-shot prompting to the LLM, enabling it to generate intermediate queries.
Inference: During the inference phase, we generate a SPARQL-PGMR query for each question using the LLM. Each URI placeholder is then replaced by its corresponding URI through the retriever, thereby forming the final SPARQL query.
Model Type F1 $\uparrow$ SQM $\uparrow$ BLEU $\uparrow$ Qid EM $\uparrow$ Pid EM $\uparrow$ URI Hallucination $\downarrow$
T5-Small FT 22.67 0.16 % 70.44 2.71 84.75 % 61.66 %
T5-Small RAG FT 67.25 64.14 % 89.85 71.98 % 94.11 % 16.01 %
T5-Small PGMR
FT 82.13 81.57 % 94.24 90.4 % 95.93 % 0.0 %
Llama 3.1 8B PEFT 22.19 0.03 % 67.33 0.53 % 65.40 % 91.04 %
Llama 3.1 8B RAG PEFT 82.35 81.24 % 94.95 89.67 % 91.10 % 7.71 %
Llama 3.1 8B PGMR
PEFT 84.47 83.33 % 94.35 92.65 % 90.80 % 0.0 %
GPT 4o 25-shot 17.27 1.45 % 55.39 4.07 % 48.92 % 85.87 %
GPT 4o RAG 25-shot 53.79 23.65 % 64.28 83.46 % 63.25 % 3.47 %
GPT 4o PGMR
25-shot 69.05 58.42 % 82.77 86.14 % 73.10 % 0.0 %
Table 2: Results on LCQUAD 2.0 (unknown URI split): Even in an out-of-distribution setting, PGMR significantly mitigates hallucinations while enhancing SQM and answer-level F1 performance over the baselines across various LLMs and settings (see Section 6).
@@ 5.1 Datasets
We evaluate our work on two widely used datasets for SPARQL query generation: LCQUAD 2.0 (Dubey et al., 2019) and QALD-10 (Usbeck et al., 2023). Even though tagged versions of these datasets exist, where the entities and relations in the question are already linked and tagged with associated URIs from the KG, we use the untagged versions. Tagging is expensive, time-consuming, and impractical in the real world. SPARQL query generation becomes a much harder problem when using untagged data (see Section 6).
LCQUAD 2.0 comprises a mix of simple and complex questions with associated SPARQL queries over Wikidata, crafted by human annotators from Amazon Mechanical Turk. The original split of LCQUAD 2.0 includes a broad dataset with 21k questions for training, 3k questions for validation, and 6k questions for testing. To specifically test how our models perform in the face of out-of-distribution URIs not seen before in training, we also construct an "unknown URI" split of LCQUAD 2.0 in line with Reyd and Zouaq (2023). Every query in the test set of this version of the dataset includes at least one URI not seen during training, thus making accurate query generation a much more complex problem. The "unknown URI" split contains 24k train, 3k validation, and 3k test questions and associated queries.
QALD-10 is a benchmarking dataset that is a part of the Question Answering over Linked Data (QALD) challenge series for KGQA using SPARQL query generation. Owing to the complexity of SPARQL queries that involve various clauses and literals, QALD-10 is recognized as one of the most challenging and practically applicable datasets in the QALD challenge series. Although each question in QALD-10 is translated into eight different languages, we only consider the English versions of the questions for this analysis. The dataset consists of 412 training questions and 394 test questions, which is much smaller than LCQUAD 2.0 but contains more complex queries. Given the limited availability of training data for QALD-10, we employ a two-step approach in our finetuning experiments: pre-training the LLM on LCQUAD 2.0 before finetuning it on QALD-10.
Model Type F1 $\uparrow$ SQM $\uparrow$ BLEU $\uparrow$ Qid EM $\uparrow$ Pid EM $\uparrow$ URI Hallucination $\downarrow$
T5-Small FT 11.30 5.07 % 34.37 11.42 % 25.38 % 66.50 %
T5-Small RAG FT 23.15 15.98 % 43.47 45.68 % 27.66 % 4.82 %
T5-Small PGMR
FT 22.12 16.75 % 44.31 42.39 % 27.41 % 0.0 %
Llama 3.1 8B PEFT 3.93 0.25 % 24.35 1.27 % 22.33 % 89.09 %
Llama 3.1 8B RAG PEFT 35.46 23.35 % 45.91 60.15 % 36.55 % 8.12 %
Llama 3.1 8B PGMR
PEFT 39.41 25.63 % 45.36 53.55 % 37.81 % 0.0 %
GPT 4o 25-shot 13.20 3.55 % 41.21 11.93 % 46.19 % 83.76 %
GPT 4o RAG 25-shot 46.52 28.68 % 51.66 74.62 % 53.55 % 11.42 %
GPT 4o PGMR
25-shot 51.19 29.19 % 54.95 65.48 % 46.19 % 0.0 %
Table 3: Results on QALD-10: PGMR consistently minimizes hallucinations across LLMs while attaining significantly better SQM and answer-level F1 scores than direct SPARQL generation, even for the more complex QALD-10 dataset, as discussed in Section 6.
@@ 5.2 Metrics
F1 Score: We execute the gold and the generated SPARQL queries and then calculate the answer-level F1 score based on the responses.
Semantic Query Match (SQM): We introduce the SQM metric, which assesses semantic equivalence rather than literal string match. Recognizing that LLMs may produce valid queries with different variable names than the reference (e.g., ?value vs. ?uri), we normalize all variables to a canonical representation (e.g., ?var0, ?var1), a technique consistent with the evaluation framework of Qi et al. (2024). Additionally, since the ordering of triples within a WHERE clause and of entities within a triple does not alter the query’s logical semantics (queries would result in the same outcome when executed), our comparison is insensitive to their permutation. Once the variables are normalized and the WHERE clause is compared in a permutation-invariant manner, the remainder of the query is subjected to an exact-match comparison.
@@
Qid-EM and Pid-EM: To assess the Qid exact match (EM) score, we directly evaluate whether the set of all entity URIs in the predicted query matches exactly with those in the reference query. Pid EM does the same for relations and properties.
@@
Refusal Accuracy is defined as the success rate with which the model identifies and correctly signals a lack of the requisite internal knowledge, thereby declining to answer the unsupported query (discussed further in Section 7).
@@ 5.3 Baselines
RAG: We additionally compare our results against the RAG paradigm, a strong baseline, wherein the question is initially processed by a retriever that retrieves $k$ URIs from the URI memory whose labels and descriptions most closely match the question text in latent space, as described in Section A.1.
PGMR: As described in Section 4, we implement the proposed PGMR, where an LLM first generates intermediate queries, followed by a retrieval step to translate them into SPARQL queries.
@@ 5.3.1 Language Models
Our evaluation incorporates a diverse range of LLMs spanning several orders of magnitude in parameter size, including models from the smaller T5 architecture (Raffel et al., 2020) (fine-tuning - FT) up to Llama 3.1 8B Instruct (parameter-efficient fine-tuning - PEFT) (Dubey et al., 2024; Mangrulkar et al., 2022) and the foundational GPT-4o (25-shot prompting) (Achiam et al., 2023).
@@ 6 Results
PGMR enhances the quality of SPARQL queries. Across all assessed LLMs, PGMR consistently enhances the accuracy and structure of SPARQL queries on both LCQUAD 2.0 and QALD-10 datasets. As seen in Tables 1, 2, and 3, PGMR improves the SQM across all baselines, LLMs, and datasets. On LCQUAD2 (original split), PGMR yields a performance gain of nearly 61.35 % over direct generation and 16.61 % over the strong RAG baseline on average across LLMs. Notably, this advantage continues on the out-of-distribution "unknown URI" split of LCQUAD2, with PGMR improving 73.89 % over direct generation and 18.09 % over RAG across LLMs, showing its robustness to shifts in data distribution. Even on the more challenging QALD-10 dataset, PGMR outperforms both direct generation and RAG across LLMs.
PGMR nearly eliminates hallucinations. From Tables 1, 2, and 3, we see that direct generation hallucinates URIs 75.42 % of the time on average across datasets and models. While RAG reduces these hallucinations to 8.71 % on average, PGMR achieves a near-complete elimination of the URI hallucination problem regardless of the LLM, dataset, and data distribution used.
Our results confirm that PGMR not only minimizes hallucinations to a great extent but also improves the accuracy and reliability of SPARQL query generation across various LLMs, datasets, and data distributions.
Figure 2: For LCQUAD2, Figure 2(a) shows the accuracy of Llama 3.1 8B PGMR on the unanswerable set as a function of the percentage of URIs removed across confidence thresholds, while Figure 2(b) shows the SQM on the answerable set (described in Section 7).
7 Analysis and discussion
Answer Refusal Experiment. A model’s ability to abstain from responding when it lacks the required information is a key determinant of its safety and trustworthiness. This mechanism directly counters model hallucination by signaling uncertainty instead of fabricating content. Such a conservative approach is indispensable in critical applications where a confidently false answer has severe consequences. We assess such robustness by randomly ablating a percentage of entity URIs from the memory, rendering dependent test queries unanswerable. We then divide the set, calculating SQM on the answerable queries and refusal accuracy on the unanswerable queries. The PGMR model architecture is inherently equipped to handle this uncertainty without additional training, thanks to the implementation of a confidence threshold applied to the retrieval similarity score. Specifically, the model executes a refusal action if the similarity score of the top-ranked retrieved URI falls below this predetermined threshold for any URI required in the generated query. Figure 2(a) demonstrates a clear positive correlation between the confidence threshold and refusal accuracy for Llama 3.1 8B PGMR. Conversely, as seen in Figure 2(b), this increase in threshold leads to a corresponding decline in SQM on the answerable subset, illustrating a direct trade-off in performance. Our analysis indicates that the optimal threshold value for balancing these two metrics is approximately 0.85. Furthermore, this confidence threshold can potentially be learned by adding a small neural network with a sigmoid head to the retriever’s encoder, but we leave this exploration for future work.
Figure 3: Llama 3.1 8B PGMR maintained robust performance on LCQUAD2, exhibiting only an approximate 4% drop in SQM despite a 9× memory size increase. This demonstrates the retriever’s resilience to memory scaling and noise (discussed in Section 7).
Memory Scaling Experiment. We investigate the scaling robustness of the PGMR framework, manipulating the entity memory size from a factor of 2 to 9 relative to the original set. This procedure involves augmenting the memory with irrelevant entities sourced from the Wikidata KG. These entities serve as distractors during the retriever’s SPARQL URI resolution. From Figure 3, we see that Llama 3.1 8B PGMR exhibits minimal performance degradation of approximately 4% on LCQUAD2, even at 9 times the memory size, validating the retriever’s robustness to expanded, noisy memory.
Inference Latency Analysis. Under equivalent constraints for maximum input and output token lengths, the inference latency of the proposed methods can be characterized by the base generation time ($T_{dg}$) and the retrieval overhead ($T_{ret}$). For a model such as Llama 3.1 8B, $T_{dg}$ is approximately 0.9 seconds per response. In contrast, the retrieval component is highly efficient, requiring only 0.0016 seconds ($T_{ret}$) and a modest 600MB of additional GPU memory. Consequently, while Direct Generation time is simply $T_{dg}$, RAG latency is defined as $T_{rag}=T_{dg}+T_{ret}$. The latency for PGMR is expressed as $T_{pgmr}=T_{dg}+(k\cdot T_{ret})$, where $k$ represents the number of URIs identified within the generated intermediate query. Given that $T_{ret}\ll T_{dg}$, the computational overhead introduced by PGMR remains negligible relative to the base model’s inference time.